OpenStudy (anonymous):

Find the sum of the following infinite geometric series, if it exists. 1/2 + (1/4) + 1/8+ (-1/16) +

OpenStudy (anonymous):

is the series 1/2 - 1/4 +1/8 -1/16?

OpenStudy (anonymous):

@sangya21 no

OpenStudy (anonymous):

r1 (1/4 & 1/2) = 1/2 r2 (1/8 & 1/4) = 1/2 r3 (-1/16 & 1/8) = -1/2 As r3 does not match with r1 and r2, thus the series does not exist